OpenStudy (miihm):

Can someone help me solve this? What is the product of these binomials? (x + 6)(x + 3) =

OpenStudy (anonymous):

\[x^{2} + 9x + 18\]

OpenStudy (miihm):

So was I supposed to add the multiply 3 and 8?

OpenStudy (miihm):

I mean 3 and 6?

OpenStudy (anonymous):

\[x ^{2}+ 3x + 6x + 18 \]

OpenStudy (anonymous):

thats what i got when i multiplied it out then i added the 3x and 6x

OpenStudy (anonymous):

You use the FOIL method to solve this. FOIL stands for First,Outside,Inside,Last and refers to the positions of the terms in the binomials. IOW, for "(x + 6)(x + 3)", you would: Multiply the (F)irst terms first: x*x = x^2 Then multiply the (O)utside terms: x*+3 = +3x Then multiply the (I)nside terms +6*x = +6x And finally you multiply the last terms: +6 * +3= +18 Now you string these products all together and you get: X^2 +3x +6x +18 Now combine common terms and you're done: X^2 +9x +18

OpenStudy (anonymous):



OpenStudy (miihm):

Ok thank you guys so mcuh! :)

OpenStudy (anonymous):

Welcome :D